# Mixed birth-death and death-birth updating in structured populations

### v1

@@ Abstract
Evolutionary graph theory (EGT) studies the effect of population structure on evolutionary dynamics. The vertices of the graph represent the $N$ individuals. The edges denote interactions for competitive replacement. Two standard update rules are death-Birth (dB) and Birth-death (Bd). Under dB, an individual is chosen uniformly at random to die, and its neighbors compete to fill the vacancy proportional to their fitness. Under Bd, an individual is chosen for reproduction proportional to fitness, and its offspring replaces a randomly chosen neighbor. Here we study mixed updating between those two scenarios. In each time step, with probability $\delta$ the update is dB and with remaining probability it is Bd. We study fixation probabilities and times as functions of $\delta$ under constant selection. Despite the fact that fixation probabilities and times can be increasing, decreasing, or non-monotonic in $\delta$, we prove nearly all unweighted undirected graphs have short fixation times and provide an efficient algorithm to estimate their fixation probabilities. We also prove that weighted directed graphs that are uniform circulations have fixation probability $1/N$ for every $\delta$. Finally, we prove exact formulas for fixation probabilities on cycles, stars, and more complex structures and classify their sensitivities to $\delta$.
@@ 1 Introduction
Evolutionary graph theory can be used to study neutral evolution [1, 2, 3], constant selection [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], or frequency dependent selection [20, 21, 22, 23, 24, 25, 26, 27]. The latter is an approach for evolutionary game theory in structured populations. In this paper, we focus on neutral evolution and constant selection.
@@
In this paper, we study $\delta$-mixed updating: in any one time step, we use death-Birth (dB) with probability $\delta$ and we use Birth-death (Bd) updating with probability $1-\delta$. There have been attempts to bridge the two worlds of dB and Bd, though those works have limited exploration into the enormous universe of the possible population structures and do not analyze fixation times [23, 13, 41]. In [23], the authors study $\delta$-mixed updating and prove that for the Prisoners’ Dilemma on a $k$-regular graph, cooperation is favored when $b/c>k/\delta$ where $b$ is the benefit from a cooperator, $c$ is the cost a cooperator pays, and $b>c>0$; if $\delta=0$, cooperation is never favored. Essentially, there is a smooth interpolation for the result of [20] between dB updating and Bd updating. In [13], the authors study graphs that maintain amplification as $\delta$ moves away from 0; they find that unlike the case in pure Bd (i.e., $\delta=0$), when $\delta>0$, amplification can only be a transient phenomenon in regards to the selection advantage. In [41], the authors use a continuous-time updating dynamics such that when normalized, there is a natural death rate $\Delta$ that triggers a death-Birth event, and there is a birth rate $1$ that triggers a Birth-death event. Then the continuous-time model is $\delta$-mixed updating with $\delta=\Delta/(1+\Delta)$. The authors find that stars on $N$ vertices with self-loops have a critical threshold at $\Delta\sim 1/\sqrt{N}$: stars are amplifiers of selection under uniform initialization when $\delta\ll 1/\sqrt{N}$ and suppressors of selection when $\delta\gg 1/\sqrt{N}$. We study fixation probability and fixation time on graphs as a function of $\delta$ in the case of neutral evolution and constant selection.
@@ 2.1 Setup
Let $G=(V,E)$ be a strongly connected directed graph on $N:=|V|$ vertices. In the case of weighted graph, there are associated edge weights $w_{uv}\in\mathbb{R}_{\geq 0}$ for each edge $u\to v\in E$. The case of an unweighted graph is modelled as all edge weights $w_{uv}$ as either: (i) zero, denoting $u\to v\not\in E$; or (ii) one, denoting $u\to v\in E$. For unweighted undirected graphs, the number of neighbors of a vertex is called its degree. The degree of vertex $u\in V$ in an unweighted undirected graph is denoted $\deg(u)$. On each vertex of $G$ lies an individual that is either a wild-type or a mutant type. Let $r>0$ be the relative fitness of a mutant and let $f_{S}\colon V\to\{1,r\}$ be the function that outputs the current fitness of an individual residing at the given vertex on the graph where $S\subseteq V$ is the set of current mutants. When the context is clear, we drop the subscript of $S$ for the mapping $f_{S}$ to simplify notation.
@@ 2.2 Dynamics
death-Birth (dB). A death-Birth (dB) step proceeds as follows. We first select an individual $v\in V$ for death uniformly at random. Then we select an incoming neighboring vertex $u\in V$ of $v$ to give birth with probability proportional to $f(u)\cdot w_{uv}$. Then the individual at location $u$ places a copy of itself at the recent vacancy of location $v$.
Birth-death (Bd). A Birth-death (Bd) step proceeds as follows. We first select an individual $u\in V$ for birth with probability proportional to its fitness, $f(u)$. Then we select an outgoing neighboring vertex $v\in V$ of $u$ to die with probability proportional to $w_{uv}$. The individual at location $u$ then places a copy of itself (and thus of its type) at the recent vacancy of location $v$.
We will refer to dB and Bd as pure updating rules since the updating rule is fixed for each step. In either case of pure updating rules, selection acts only on the birthing individual. Next, we define an interpolation of the pure updating rules.
Mixed $\delta$-updating. Let $\delta\in[0,1]$ be the probability a given time step will be a dB step. With remaining probability, the step will be a Bd step. Note that we recover pure dB updating when $\delta=1$ and recover pure Bd updating when $\delta=0$. When $\delta<1/2$, the process is Bd-biased; when $\delta>1/2$, the process is dB-biased. When $\delta=1/2$, we refer to the process as unbiased.
@@ 2.3 Fixation probability and time
For $S\subseteq V$, let $w(S)=\sum_{u\in V}f_{S}(u)=N+(r-1)|S|$ be the total fitness of the population given that mutants reside at the locations in $S$; also, let $w_{u}(S)=\sum_{v:(u,v)\in E}f_{S}(v)$ be the total fitness of the neighbors of $u\in V$. Eventually, the population becomes entirely mutant—we call this fixation—or entirely wild-type—we call this extinction. In either case, we say that absorption has occurred. We denote the probability of fixation given the mutants start at $S_{0}\subseteq V$ as $\mathsf{fp}_{r}^{\delta}(G,S_{0})$. We also denote the uniform initialization fixation probability as $\mathsf{fp}_{r}^{\delta}(G):=\frac{1}{N}\sum_{u\in V}\mathsf{fp}_{r}^{\delta}(G,\{u\})$. When dealing with times, we focus solely on the expected time rather than the entire distribution. For an initial mutant set $S_{0}\subseteq V$, we denote the (expected) time until fixation, extinction, and absorption as $\mathsf{FT}_{r}^{\delta}(G,S_{0})$, $\mathsf{ET}_{r}^{\delta}(G,S_{0})$, and $\mathsf{AT}_{r}^{\delta}(G,S_{0})$, respectively. Since we are interested in upper bounds for times, we drop the mutant initial set from the notation for time to denote the maximum time over all possible initial mutant sets. Thus, for example, $\mathsf{FT}_{r}^{\delta}(G):=\max_{S_{0}\subseteq V}\mathsf{FT}_{r}^{\delta}(G,S_{0})$. In the notations for probabilities and times, we use $\delta=\textrm{dB}$ to mean $\delta=1$ and $\delta=\textrm{Bd}$ to mean $\delta=0$.
@@ 3 Results
Figure 2: Fixation probabilities of various graphs under mixed $\delta$-updating in neutral evolution for $N=7$. In both figures, the dashed line at $1/N$ is the fixation probability of the Moran process (well-mixed) in neutral evolution. For each graph depicted, the small star symbol on the differently shaded vertex indicates the initial mutant location. a, i) A star graph with the initial mutant location at the center. Relative to $\rho$ under the unbiased process, Bd-biased processes have a lower $\rho$ but dB-biased processes have a higher $\rho$. a, ii) A star graph with the initial mutant location at a leaf. Relative to $\rho$ under the unbiased process, Bd-biased processes have a higher $\rho$ but dB-biased processes have a lower $\rho$. a, iii) A graph composed of a star and two additional edges. Relative to $\rho$ under the unbiased process, both pure Bd and pure dB processes have a lower $\rho$. However, panel b, and iv) shows that the graph from a, iii) has higher $\rho$ than the unbiased process when $\delta$ is slightly smaller than $1/2$. This example shows that $\rho$ can be non-monotonic in $\delta$. Further, it is not possible that both pure Bd and pure dB have higher $\rho$ than the unbiased process (see Theorem A.1).
3.1 Neutral evolution fixation probability
@@
Suppose that $G=(V,E)$ is an unweighted undirected graph and the initial set of mutants reside at $S_{0}\subseteq V$. We first show by example that $\mathsf{fp}_{r=1}^{\delta}(G,S_{0})$ as a function of $\delta$ can be increasing, decreasing, or non-monotone (see Figure 2). For the pure Bd process, it is known that $\mathsf{fp}_{r=1}^{\mathrm{Bd}}(G,S_{0})\propto\sum_{u\in S_{0}}1/\deg(u)$ [42]; for the pure dB process, it is known that $\mathsf{fp}_{r=1}^{\mathrm{dB}}(G,S_{0})\propto\sum_{u\in S_{0}}\deg(u)$ [40]. Secondly, we prove that for general $\delta$ and neutral evolution, the fixation probability is additive in the initial mutant set like it is in the pure Bd and dB processes; that is $\mathsf{fp}_{r=1}^{\delta}(G,S_{0})=\sum_{u\in S_{0}}\mathsf{fp}_{r=1}^{\delta}(G,\{u\})$ (see Lemma B.1). Finally, we prove the following result for $\delta=1/2$.
Theorem 3.1 (Neutral evolution fixation probability).
Let $G=(V,E)$ be a undirected unweighted graph. Suppose the initial mutants reside at $S_{0}\subseteq V$. Then $\mathsf{fp}_{r=1}^{\delta=1/2}(G,S_{0})=|S_{0}|/N$.
3.2 Uniform circulations
Figure 3: A uniform circulation on $N=7$ vertices with $m=1$. Thus each vertex has a total incoming sum of $1$ and a total outgoing sum of $1$. The weight of each directed edge is labeled near the edge arrowhead. The thickness of each edge is proportional to its weight. Self-loops have no arrow depicted. The vertex highlighted on the top has incoming weight $1/3+1/6+1/4+1/4=1$ and outgoing weight $1/6+1/3+1/4+1/4=1$. The vertex highlighted on the bottom has incoming weight $1/4+1/6+1/4+1/3=1$ and outgoing weight $1/2+1/6+1/3=1$.
Suppose we have a directed weighted graph $G=(V,E)$ on $N$ vertices. One can ask: which graphs have the same fixation probability as a complete graph, $K_{N}$? It is known for the Bd process that if $G$ has the property that the probability that death occurs at a particular vertex is the same for all other vertices as well—known as an isothermal graph—then $\mathsf{fp}_{r}^{\mathrm{Bd}}(G)=\mathsf{fp}_{r}^{\mathrm{Bd}}(K_{N})$ [30]. For the pure dB process, the isothermal theorem does not hold [38]; furthermore it is even the case that $\mathsf{fp}_{r}^{\mathrm{Bd}}(K_{N})\neq\mathsf{fp}_{r}^{\mathrm{dB}}(K_{N})$ in general [6]. We prove that in neutral evolution, graphs where the outgoing and incoming weights sum to the same universal quantity for all vertices—we call this a uniform circulation—have fixation probability the same as a complete graph, regardless of $\delta$.
Definition 3.1.
A weighted, directed graph $G=(V,E)$ is called a uniform circulation if there exists some $m>0$ such that $\sum_{v\in V}w_{uv}=\sum_{v\in V}w_{vu}=m$ for all $u\in V$.
A graph is isothermal if it is a uniform circulation, but the converse is not generally true. Thus, the uniform circulation property is a more stringent condition for a graph than isothermality. When the directed graph is unweighted, uniform circulation means that the graph is both regular and Eulerian. In the case of an unweighted undirected graph, uniform circulation means that the graph is regular.
Theorem 3.2 (Uniform circulation theorem).
Let $G=(V,E)$ be a weighted strongly connected directed graph. Suppose the initial mutants reside at $S_{0}\subseteq V$. If $G$ is a uniform circulation then $\mathsf{fp}_{r=1}^{\delta}(G,S_{0})=|S_{0}|/N$, regardless of $\delta$.
One could similarly ask: for fixed $\delta$, which graph families $(G_{N}^{\delta})_{N\geq 1}$ have the property that $\mathsf{fp}_{r}^{\delta}(G_{N}^{\delta})=\mathsf{fp}_{r}^{\delta}(K_{N})$ for all $r$ and $N$? We leave this question open.
3.3 Fixation and absorption times
For undirected unweighted graph families $(G_{N})_{N\geq 1}$, it is known that for the pure Bd process, $\mathsf{AT}_{r}^{\textrm{Bd}}(G_{N})$ is $o(N^{3+\varepsilon})$ for all $\varepsilon>0$ and $r\neq 1$ [43, 44]. For $r=1$, it is known $\mathsf{AT}_{r=1}^{\textrm{Bd}}(G_{N})$ is $O(N^{6})$ [1]. For pure dB, less is known. Generally, we know that for $r=1$ we have $\mathsf{AT}_{r=1}^{\textrm{dB}}(G_{N})$ is $O(N^{5})$ [45]. Relating fixation times to absorption times, the law of total expectation on the absorption time gives the bound $\mathsf{FT}_{r}^{\delta}(G,S_{0})\leq\mathsf{AT}_{r}^{\delta}(G,S_{0})/\mathsf{fp}_{r}^{\delta}(G,S_{0})$ for any $S_{0}\subseteq V$, $r>0$, and $\delta\in[0,1]$. For any undirected unweighted graph $G=(V,E)$ on $N$ vertices when $r>1$ and $\delta=0,1$, we have $\mathsf{fp}_{r}^{\delta}(G)\geq\mathsf{fp}_{r=1}^{\delta}(G)$ where the right hand side of the inequality is $\Omega(N^{-2})$ [^1] For functions $f$ and $g$ that map $\mathbb{N}$ to $\mathbb{R}_{>0}$, we say that “$f(n)$ is $\Omega(g(n))$” if there exist some constants $c\in\mathbb{R}_{>0}$ and $n_{0}\in\mathbb{N}$ such that $f(n)\geq c\cdot g(n)$ for every $n\geq n_{0}$.. Thus for $r>1$ we have $\mathsf{FT}_{r}^{\mathrm{Bd}}(G_{N})$ is $o(N^{5+\varepsilon})$ for all $\varepsilon>0$. Similar to fixation probabilities, fixation times in the mixed $\delta$-updating process can exhibit a large range of behaviors relative to $\delta$ (see Figure 4).
Figure 4: a, Fixation times of various graphs under mixed $\delta$-updating in neutral evolution for $N=7$. For each graph depicted, the small star symbol on the differently shaded vertex indicates the initial mutant location. i) The fixation time increases and then decreases relative to $\delta$. ii) The fixation time decreases and then increases relative to $\delta$. iii) The fixation time always increases relative to $\delta$. iv) The fixation time always decreases relative to $\delta$.
We prove two results related to the fixation and absorption times. We state the first result.
Theorem 3.3.
For all unweighted undirected graphs $G=(V,E)$ on $N$ vertices, we have $\mathsf{FT}_{r=1}^{\delta=1/2}(G)\leq\frac{r}{r-1}\cdot N^{5}$.
To motivate our result on fixation time, we introduce some notation. Let $\mathscr{G}_{\leq N}$ be the set of all $N$-length finite sequences of unweighted undirected graphs $(G_{1},\ldots,G_{N})$ such that graph $G_{k}$ has $k$ vertices. For any subset $\mathscr{H}_{\leq N}\subseteq\mathscr{G}_{\leq N}$, denote both
@@ 3.3 Fixation and absorption times
    $\mathscr{H}:=\lim\limits_{N\to\infty}\mathscr{H}_{\leq N}$, and

### v2

# Mixed updating in structured populations

@@ Abstract
Evolutionary graph theory (EGT) studies the effect of population structure on evolutionary dynamics. The vertices of the graph represent the $N$ individuals. The edges denote interactions for competitive replacement. Two standard update rules are death-Birth (dB) and Birth-death (Bd). Under dB, an individual is chosen uniformly at random to die, and its neighbors compete to fill the vacancy proportional to their fitness. Under Bd, an individual is chosen for reproduction proportional to fitness, and its offspring replaces a randomly chosen neighbor. Here we study mixed updating between those two scenarios. In each time step, with probability $\delta$ the update is dB and with remaining probability it is Bd. We study fixation probabilities and times as functions of $\delta$ under neutral evolution and constant selection. Despite the fact that fixation probabilities and times can be increasing, decreasing, or non-monotonic in $\delta$, we prove nearly all unweighted undirected graphs have short fixation times and provide an efficient algorithm to estimate their fixation probabilities. Finally, we prove exact formulas for fixation probabilities on cycles, stars, and more complex structures and classify their sensitivities to $\delta$.
@@ 1 Introduction
Evolutionary graph theory can be used to study neutral evolution [1, 2, 3], constant selection [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], or frequency dependent selection [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. The latter is an approach for evolutionary game theory in structured populations. In this paper, we focus on neutral evolution and constant selection.
@@
In this paper, we study $\delta$-mixed updating: in any one time step, we use death-Birth (dB) with probability $\delta$ and we use Birth-death (Bd) updating with probability $1-\delta$. There have been attempts to bridge the two worlds of dB and Bd, though those works have limited exploration into the enormous universe of the possible population structures and do not analyze fixation times [26, 13, 45]. In [26], the authors study $\delta$-mixed updating and prove that for the Prisoners’ Dilemma on a $k$-regular graph, cooperation is favored when $b/c>k/\delta$ where $b$ is the benefit from a cooperator, $c$ is the cost a cooperator pays, and $b>c>0$; if $\delta=0$, cooperation is never favored. Essentially, there is a smooth interpolation for the result of [46] between dB updating and Bd updating. In [13], the authors study graphs that maintain amplification as $\delta$ moves away from 0; they find that unlike the case in pure Bd (i.e., $\delta=0$), when $\delta>0$, amplification can only be a transient phenomenon in regards to the selection advantage. In [45], the authors use a continuous-time updating dynamics such that when normalized, there is a natural death rate $\Delta$ that triggers a death-Birth event, and there is a birth rate $\varphi$ that triggers a Birth-death event where $\varphi$ is the average fitness of the current population. Then their continuous-time model is $\delta$-mixed updating with $\delta=\Delta/(\varphi+\Delta)$. The authors find that stars on $N$ vertices with self-loops have a critical threshold at $\Delta\sim 1/\sqrt{N}$: stars are amplifiers of selection under uniform initialization when $\delta\ll 1/\sqrt{N}$ and suppressors of selection when $\delta\gg 1/\sqrt{N}$. We study fixation probability and fixation time on unweighted undirected graphs as a function of $\delta$ in the case of neutral evolution and constant selection.
@@ 2.1 Setup
Let $G=(V,E)$ be a connected undirected graph on $N:=|V|$ vertices. The number of neighbors of a vertex in a graph is called its degree. The degree of vertex $u\in V$ is denoted $\deg(u)$. On each vertex of $G$ lies an individual that is either a wild type or a mutant type. Mutants have relative fitness $r\geq 1$ and wild-types have fitness $1$. When $r=1$, the evolution is neutral; when $r>1$, mutants have a selective advantage (constant selection).
@@ 2.2 Dynamics
death-Birth (dB). A death-Birth (dB) step proceeds as follows. We first select an individual $v\in V$ for death uniformly at random. Then we select an incoming neighboring vertex $u\in V$ of $v$ to give birth proportional to fitness. Then the individual at location $u$ places a copy of itself at the recent vacancy of location $v$.
Birth-death (Bd). A Birth-death (Bd) step proceeds as follows. We first select an individual $u\in V$ for birth proportional to fitness. Then we select an outgoing neighboring vertex $v\in V$ of $u$ to die uniformly at random. The individual at location $u$ then places a copy of itself (and thus of its type) at the recent vacancy of location $v$.
We will refer to dB and Bd as pure updating rules since the updating rule is fixed for each step. Next, we define an interpolation of the pure updating rules.
Mixed $\delta$-updating. Let $\delta\in[0,1]$ be the probability a given time step will be a dB step. With remaining probability, the step will be a Bd step. We only consider the case when the value of $\delta$ is constant throughout the entire process. Note that we recover pure dB updating when $\delta=1$ and recover pure Bd updating when $\delta=0$. When $\delta<1/2$, the process is Bd-biased; when $\delta>1/2$, the process is dB-biased. When $\delta=1/2$, we refer to the process as unbiased. See Figure 1 for an illustration of each step of the process.
@@ 2.3 Fixation probability and time
Eventually, the population becomes entirely mutant—we call this fixation—or entirely wild-type—we call this extinction. In either case, we say that absorption has occurred. We denote the probability of fixation given mutant fitness $r\geq 1$ and initial set $S_{0}\subseteq V$ as $\mathsf{fp}_{r}^{\delta}(G,S_{0})$. We also denote the uniform initialization fixation probability as $\mathsf{fp}_{r}^{\delta}(G):=\frac{1}{N}\sum_{u\in V}\mathsf{fp}_{r}^{\delta}(G,\{u\})$. When dealing with times, we focus solely on the expected time rather than the entire distribution. For an initial mutant set $S_{0}\subseteq V$, we denote the (expected) time until fixation, extinction, and absorption as $\mathsf{FT}_{r}^{\delta}(G,S_{0})$, $\mathsf{ET}_{r}^{\delta}(G,S_{0})$, and $\mathsf{AT}_{r}^{\delta}(G,S_{0})$, respectively. Since we are interested in upper bounds for times, we drop the mutant initial set from the notation for time to denote the maximum time over all possible initial mutant sets. Thus, for example, $\mathsf{FT}_{r}^{\delta}(G):=\max_{S_{0}\subseteq V}\mathsf{FT}_{r}^{\delta}(G,S_{0})$. In the notations for probabilities and times, we use $\delta=\textrm{dB}$ to mean $\delta=1$ and $\delta=\textrm{Bd}$ to mean $\delta=0$. When the fitness is neutral ($r=1$), we write $\mathsf{fp}_{r=1}^{\delta}$, $\mathsf{FT}_{r=1}^{\delta}$, etc.
@@ 3 Results
Figure 2: Fixation probabilities, $\rho$, of various graphs under mixed $\delta$-updating for $N=7$ in neutral evolution. In both figures, the dashed line at $1/N$ is the fixation probability of the Moran process (well-mixed) in neutral evolution. For each graph depicted, the small star symbol on the differently shaded vertex indicates the initial mutant location. a, i) A star graph with the initial mutant location at the center. Relative to $\rho$ under the unbiased process, Bd-biased processes have a lower $\rho$ but dB-biased processes have a higher $\rho$. a, ii) A star graph with the initial mutant location at a leaf. Relative to $\rho$ under the unbiased process, Bd-biased processes have a higher $\rho$ but dB-biased processes have a lower $\rho$. a, iii) A graph composed of a star and two additional edges. Relative to $\rho$ under the unbiased process, both pure Bd and pure dB processes have a lower $\rho$. However, panel b, and iv) shows that the graph from a, iii) has higher $\rho$ than the unbiased process when $\delta$ is slightly smaller than $1/2$. This example shows that $\rho$ can be non-monotonic in $\delta$. Further, it is not possible that both pure Bd and pure dB have higher $\rho$ than the unbiased process (see Lemma D.1).
3.1 Neutral evolution
@@
Suppose that $G=(V,E)$ is graph and the initial set of mutants reside at $S_{0}\subseteq V$. For the pure Bd process in neutral evolution, it is known that $\mathsf{fp}_{r=1}^{\mathrm{Bd}}(G,S_{0})\propto\sum_{u\in S_{0}}1/\deg(u)$ [47]; for the pure dB process, it is known that $\mathsf{fp}_{r=1}^{\mathrm{dB}}(G,S_{0})\propto\sum_{u\in S_{0}}\deg(u)$ [44]. Due to the additivity of the fixation probability in the pure processes, $\mathsf{fp}_{r=1}^{\mathrm{Bd}}(G)=\mathsf{fp}_{r=1}^{\mathrm{dB}}(G)=1/N$.
First, we prove that under general $\delta$, the fixation probability is similarly additive in the initial mutant set like it is in the pure Bd and dB processes; that is $\mathsf{fp}_{r=1}^{\delta}(G,S_{0})=\sum_{u\in S_{0}}\mathsf{fp}_{r=1}^{\delta}(G,\{u\})$ (see Lemma B.1). Since $\mathsf{fp}_{r=1}^{\delta}(G)=(1/N)\cdot\sum_{u\in V}\mathsf{fp}_{r=1}^{\delta}(G,\{u\})$, additivity easily yields the following result.
Theorem 3.1 (Uniform initialization).
Averaging uniformly over all initial conditions, the fixation probability of a neutral mutant is $1/N$ on any graph for any value of delta.
In our notation, Theorem 3.1 means that $\mathsf{fp}_{r=1}^{\delta}(G)=1/N$ for any graph $G$ and any value of $\delta$. Thus mixed $\delta$-updating under uniform initialization is well-behaved. Let $u\in V$. It is known that if one of $\mathsf{fp}_{r=1}^{\mathrm{Bd}}(G,\{u\})$ or $\mathsf{fp}_{r=1}^{\mathrm{dB}}(G,\{u\})$ is larger than $1/N$, then the other is at most $1/N$ (see Lemma D.1). Surprisingly, we prove that no matter the fixation probabilities of the pure processes with initial mutant location $u$, the unbiased process always has fixation probability $1/N$ when the initial mutant location is $u$. This means that even if both pure processes have fixation probability less than $1/N$ for initial mutant location $u$, the unbiased process has a higher fixation probability than both the pure processes for initial mutant location $u$.
Theorem 3.2 (Arbitrary initialization).
If $\delta=1/2$ (which is the unbiased case) then the fixation probability of a neutral mutant starting in an arbitrary location is $1/N$ on any graph.
In our notation, Theorem 3.2 states that for $G=(V,E)$ and $u\in V$ we have $\mathsf{fp}_{r=1}^{\delta=1/2}(G,\{u\})=1/N$. As demonstrated by Figure 2, the quantity $\mathsf{fp}_{r=1}^{\delta}(G,\{u\})$ as a function of $\delta$ can be increasing, decreasing, concave, convex, or non-monotone. Despite this possible strange behavior, we prove that the fixation probabilities on regular graphs are not affected by $\delta$.
Theorem 3.3 (Fixation probability of regular graphs).
The fixation probability of a neutral mutant starting in an arbitrary location on a regular graph is $1/N$ for any value of $\delta$.
In fact, we prove something stronger: there is a broad class of graphs whose fixation probabilities are monotone in $\delta$. Such graphs—called bidegreed—have a vertex cut such that the vertices in the cut all have the same degree and the vertices outside the cut all have the same degree.
Definition 3.1 (Bidegreed graph).
A bidegreed graph is an undirected graph with the following property: for each vertex the number of neighbors is either $d_{1}$ or $d_{2}$.
If $d_{1}=d_{2}$ then the graph is regular. The following statements are proven in Appendix E.1.
Theorem 3.4 (Bidegreed graphs).
On a bidegreed graph with $d_{1}<d_{2}$, the fixation probability of a neutral mutant starting on a vertex with degree $d_{1}$ is a strictly decreasing function of $\delta$. The fixation probability of a neutral mutant starting on a vertex with degree $d_{2}$ is a strictly increasing function of $\delta$.
Let $n_{i}$ denote the number of vertices in a bidegreed graph $G$ with degree $d_{i}$ for $i=1,2$. We denote $\rho_{i}(\delta)$ as $\mathsf{fp}_{r=1}^{\delta}(G,\{u\})$ where $u$ is a vertex with degree $d_{i}$. The sensitivity of a vertex with degree $d_{i}$ is $s_{i}(\delta):=|\rho^{\prime}_{i}(\delta)|$. In addition to Theorem 3.4, we prove that for $i=1,2$,
@@
    if $n_{1}>n_{2}$, $s_{i}$ has its maximum at $\delta=1$ and minimum at $\delta=0$;